Problem 10.33

Complete the following table for an ideal gas:
P V n T
2.10 atm 1.20 L 0.520 mol ? K
0.320 atm 0.270 L ? mol 32 ∘C
660 torr ? L 0.335 mol 370 K
? atm 595 mL 0.270 mol 280 K

Part A

Complete the first row of the table.

T = 59.1   K  

SubmitMy AnswersGive Up

Correct

Significant Figures Feedback: Your answer 59.02 K was either rounded differently or used a different number of significant figures than required for this part.

Part B

Complete the second row of the table.

n =

.0027

  mol  

SubmitMy AnswersGive Up

Incorrect; Try Again; 5 attempts remaining

Part C

Complete the third row of the table.

V =

.1878

  L  

SubmitMy AnswersGive Up

Incorrect; Try Again; 5 attempts remaining

Part D

Complete the fourth row of the table.

P =

78.7

  atm  

SubmitMy AnswersGive Up

Incorrect; Try Again; 5 attempts remaining
Answer #1

A)

Given:

P = 2.1 atm

V = 1.2 L

n = 0.520 mol

use:

P * V = n*R*T

2.1 atm * 1.2 L = 0.52 mol* 0.0821 atm.L/mol.K * T

T = 59.0 K

B)

Given:

P = 0.32 atm

V = 0.27 L

T = 32.0 oC

= (32.0+273) K

= 305 K

find number of moles using:

P * V = n*R*T

0.32 atm * 0.27 L = n * 0.08206 atm.L/mol.K * 305 K

n = 3.45*10^-3 mol

= 0.00345 mol

C)

Given:

P = 660.0 torr

= (660.0/760) atm

= 0.868421 atm

n = 0.335 mol

T = 370.0 K

use:

P * V = n*R*T

0.868421 atm * V = 0.335 mol* 0.0821 atm.L/mol.K * 370 K

V = 11.7 L

D)

Given:

V = 595.0 mL

= (595.0/1000) L

= 0.595 L

n = 0.270 mol

T = 280.0 K

use:

P * V = n*R*T

P * 0.595 L = 0.27 mol* 0.0821 atm.L/mol.K * 280 K

P = 10.4 atm

Answer: 10.4 atm